OpenStudy (anonymous):

2sin^2x + 2cosx - 1 =0 PLEASE HELP

OpenStudy (solomonzelman):

\[2\sin^2x + 2cosx - 1 =0 \]\[2(1-\cos^2x) + 2cosx - 1 =0\]\[2-2\cos^2x + 2cosx - 1 =0\]\[1-2\cos^2x + 2cosx =0\] let cos(x)=a can you do it from here ?